# MASQUE: Localized Adversarial Makeup Generation with Text-Guided Diffusion Models

### v2

@@ 4.2 Detailed Design
To address the limitations of existing AFR methods discussed in Section 4.1, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced. Figure 2 illustrates the pipeline of MASQUE with its pseudocode detailed in Algorithm 1 in Appendix A. The makeup generation process is guided by a user-defined text prompt , refined via cross-attention fusion with a mask. Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR models without compromising visual fidelity. In particular, our method leverages the state-of-the-art Stable Diffusion (Rombach et al., 2022) adapted to generate high-quality adversarial makeups by iterative denoising a random noise latent conditioned on a text embedding . To be more specific, Stable Diffusion is trained to predict the added noise via:
@@
Enhancing Semantic Edits and Locality. To ensure precise localization, we generate a mask that defines the region for modification. To achieve this, we embed the prompt using a Sentence Transformer (Reimers & Gurevych, 2019) model and compare it to embeddings of predefined facial regions. The closest match determines the relevant area for the edit. For instance, if the prompt specifies “a face with red lipstick”, the model identifies lips as the target and generates a lip-area mask. This ensures that our adversarial makeup perturbations remain focused on the correct facial region and do not unintentionally alter other parts of the face. Once the target region is determined, we enhance the influence of makeup-related tokens by maximizing their attention within . This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality (Mao et al., 2023). Specifically, we optimize:
@@
Pariwise Adversarial Guidance. Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios. By contrast, our approach leverages a pair of face images from the same individual, where serves as the guide image. This strategy highlights a significant issue with naively using the distance from the original image as the adversarial loss. Note that in the standard diffusion denoising process, the objective is to generate images similar to the original. If we merely maximize the distance from the original image as an adversarial loss, it can potentially create conflicting objectives, leading to an unstable performance in both image quality and adversarial effectiveness. In our evaluations, we test the impact of the number of guide images on MASQUE’s performance (Appendix C.1) and whether using self-augmentation without a separate guide image () in MASQUE can be an alternative (Appendix C.2).
Adversarial and Image Quality Trade-off. We introduce adversarial perturbations during the later stage of the diffusion process, ensuring the coarse structure remains intact while subtly altering identity-specific features. To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure:
@@ 5 Experiments
In this section, we conduct experiments to benchmark the performance of MASQUE in dodging success and visual aspects, with comparisons to state-of-the-art makeup-based AFR methods. Due to space limits, we present additional experiments, ablation studies, and visualizations in Appendices B-E.
@@ 5.1 Experimental Setup
Dataset. To reflect the high-quality nature of online face images, we use high-resolution images () from each of the CelebA-HQ (Karras et al., 2017) and VGG-Face2-HQ (Chen et al., 2024) datasets. For each dataset, we randomly sample 100 identities, each with three images: a probe (to be protected), a reference (as the guide image), and a gallery image (stored in the attacker’s facial recognition model). Note that the VGG-Face2-HQ dataset contains face images under challenging conditions, such as pose variation and natural occlusions like sunglasses and hair.
Configuration. We compare MASQUE against the following baselines: TIP-IM (Yang et al., 2021), AMT-GAN (Hu et al., 2022), C2P (Shamshad et al., 2023), and DiffAM (Sun et al., 2024). For C2P, both its impersonating version, C2P (I), and its dodging version, C2P (D), are assessed. We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 (Deng et al., 2019), IRSE50 (Hu et al., 2018), FaceNet (Schroff et al., 2015), and MobileFace (Chen et al., 2018). Without explicitly mentioning, we set in MASQUE (using a single separate guide image of the same individual). For reference-based methods, we selected images from the makeup dataset (Li et al., 2018) used during their pre-training that best matched the given prompt. In addition to public FR models, we evaluate AFR methods against two commercial APIs: Face++111https://www.faceplusplus.com/face-comparing/ and Luxand222https://luxand.cloud/face-api.
Evaluation Metric. We employ DSR to evaluate the protection effectiveness of AFR methods. We measure visual quality using LPIPS, PSNR, and SSIM (Wang et al., 2004). Lower LPIPS and higher PSNR/SSIM indicate better preservation in terms of perceptual and structural similarity. Besides, we assess prompt following ability using metrics like concentration and hue accuracy. In Appendix B, we provide detailed descriptions of the hyperparameters used in MASQUE, implementation details for external image references, and formal definitions of the evaluation metrics.
@@ 5.2 Main Results
Dodging Success. Table 2 presents the DSR for both face verification and rank-1 face identification in black-box scenarios, evaluated using four widely used pre-trained FR feature extractors (Deng et al., 2019; Hu et al., 2018; Schroff et al., 2015; Chen et al., 2018). For each target model, the remaining ones serve as surrogate models to simulate the black-box scenario, with results averaged across three makeup styles (“red lipstick”, “blue eyebrow”, and “pink eyeshadow”). Table 2 demonstrates that MASQUE () consistently outperforms existing baselines across all the tested configurations, achieving an average DSR of for identification and for verification. Besides, even using no guide image, MASQUE still outperforms TIP-IM, the second-best AFR method, by a large margin, confirming its strong protection performance under dodging scenarios. Besides, we test our method against two commercial FR APIs, Face++ and Luxand, in verification mode, which assigns similarity scores from to . Figure 3 shows the results. As a proprietary model with unknown training data and parameters, it serves as a realistic testbed for evaluating the effectiveness of AFR methods. Our approach achieves the lowest similarity score across both APIs, demonstrating effectiveness in both open-source and closed-source settings, reinforcing its real-world applicability.
| TIP-IM | AMT-GAN | C2P(I) | C2P(D) | DiffAM | MASQUE () | MASQUE () | |
| LPIPS() | |||||||
| PSNR() | |||||||
@@
Visual Quality. MASQUE achieves superior image quality across multiple evaluation metrics, as summarized in Table 3. While TIP-IM attains the highest PSNR and SSIM due to its small perturbation constraint, these pixel-level metrics often fail to reflect perceptual quality (see Figure 1 for qualitative comparison results). In contrast, our approach prioritizes perceptual consistency, balancing content fidelity and visual realism, as demonstrated by its strong LPIPS performance.
@@
Moreover, we examine the localization capability and prompt adherence of MASQUE based on spatial precision and color accuracy. For spatial adherence, we use concentration, which quantifies the proportion of modifications within a binary mask of the target area. Higher scores indicate better localization, while lower scores suggest spillover. For color precision, we evaluate hue accuracy by extracting modified pixels within the mask, converting them to HSV space, and measuring how many fall within the expected hue range. Figure 5 shows that MASQUE consistently achieves the highest concentration scores and competitive hue accuracy across prompts, ensuring precise localization with accurate color application. In contrast, DiffAM struggles with blue eyebrows, likely due to its framework restricting makeup transfer to specific facial regions, limiting its flexibility. In Appendix D.1, we conduct additional experiments using other distance metrics with respect to both in- and out-mask regions, illustrating the strong localization capability of MASQUE. These results confirm that MASQUE achieves strong prompt adherence by maintaining spatial precision while accurately reflecting the intended color attributes.
6 Further Analysis
@@
Controllable Makeup Generation. Figure 4 illustrates the strong controllability of MASQUE over text makeup prompts across three dimensions: sequential editing, intensity control, and style control. Makeup can be applied progressively to multiple facial regions, with each layer enhancing protection, as shown by decreasing similarity scores. Intensity is adjustable via text prompts (e.g., “vivid”), affecting appearance with minimal impact on effectiveness. Style prompts range from natural (e.g., “tanned skin”) to extreme (e.g., “clown makeup”), with broader coverage offering stronger protection (see Appendix E for additional visualizations of MASQUE-generated images).
| Method | Identification | Verification | ||||||
| MASQUE | ||||||||
| + Resize | ||||||||
| + Compression | ||||||||
| + Gaussian Noise | ||||||||
| + Blur | ||||||||
| + Fog | ||||||||
| + Makeup Removal | ||||||||
| + DDPM Denoising | ||||||||
@@
Robustness to Image Transformation. We evaluate the robustness of MASQUE using the same CelebA-HQ images and four FR models used in the main experiments, with the prompt “red lipstick”. We then subject the edited images to a variety of transformations, standard post-processing such as JPEG compression, blurring, natural corruptions like ImageNet-C, and adaptive attacks such as DiffAM’s makeup removal and DDPM-based denoising. As shown in Table 4, MASQUE maintains high dodging success rates across all cases, indicating strong resilience to real-world degradations and adaptive threats. This robustness may arise from the invariance of FR models, which are trained to withstand such transformations, allowing MASQUE’s protection to persist. Also, we test the performance of MASQUE against adversarial purification (Nie et al., 2022), where the results are presented and discussed in Appendix D.4.
Computational Analysis. MASQUE does not require fine-tuning, operating directly with a pretrained Stable Diffusion model. In contrast, Clip2Protect fine-tunes per image (~s), and DiffAM performs per-dataset fine-tuning (~min) for each style or identity. At inference, MASQUE runs a one-time iterative optimization (~s), slightly slower than Clip2Protect (~s) and DiffAM (~s). Despite this, MASQUE offers key advantages: it supports arbitrary prompts and identities without retraining, making it more flexible and general. Under typical AFR scenarios, where protection is applied once before sharing, this adaptability outweighs the modest runtime difference.
@@ 7 Conclusion
We introduced MASQUE, a diffusion-based anti-facial recognition framework that applies adversarial makeup via text prompts, enabling localized modifications with automatically generated masks. Unlike prior AFR methods, MASQUE achieves high dodging success rates while preserving visual quality without relying on external IDs, mitigating the potential risks of targeted misuse. Our experiments demonstrated the superiority of MASQUE over existing baselines, where we highlight the important role of pairwise adversarial guidance. Future work may focus on optimizing the efficiency for real-time applications and studying the robustness against adaptively evolving FR models (Fan et al., 2025) to further strengthen MASQUE’s efficacy for facial privacy protection.
Availability
The code implementations of our method and for reproducing our experimental results are open-sourced and available at: https://github.com/TrustMLRG/MASQUE.
Appendix A Algorithm Pseudocode
@@ Appendix B Detailed Experimental Settings
| #Guide Images | Iden. DSR | Veri. DSR | Confidence | Similarity | LPIPS |
@@ C.4 Mask and Perturbation Localization
| TIP-IM | AMT-GAN | C2P(I) | C2P(D) | DiffAM | MASQUE () | |
| LPIPS() | ||||||
| PSNR() | ||||||
@@ D.1 Localized Editing
To evaluate whether edits are confined to the desired region, images are divided into in-mask and out-mask regions using a binary mask. The corresponding areas in both images are isolated via element-wise multiplication. The areas are then used to normalize similarity metrics proportionally, ensuring fair comparisons. Specifically, we employ DISTS, LPIPS, and PieAPP as evaluation metrics in this experiment: DISTS measures perceptual dissimilarity based on structure and texture, LPIPS uses deep features, and PieAPP reflects human perceptual preferences.
Table 7 compares the effectiveness of localized edits, where higher values indicate stronger localization of perturbations. TIP-IM (Yang et al., 2021), as a noise-based method, applies pixel-wise adversarial perturbations uniformly, resulting in the smallest difference for all metrics due to minimal distinction between in-mask and out-mask regions. In contrast, our method introduces significant perturbations in the in-mask region, leading to poorer metrics there, but achieves the best or second-best results in the out-mask region, indicating minimal disruption to untouched areas.
@@ D.4 Robustness against Adversarial Purification
D.5 Limitations and Broader Impact
@@
While our method is designed to defend against unauthorized or invasive FR systems, it may raise dual-use concerns. Specifically, the same techniques could potentially be misused to evade legitimate or authorized FR systems, such as those used in security or law enforcement contexts. We believe it is important to anticipate and acknowledge these risks, and we encourage future work to consider safeguards or contextual constraints that limit misuse while preserving individual privacy protections.
@@ Appendix E Additional Visualizations
Additional visual results of image-guided makeup transfer are provided in Figure 8 and Figure 9 to accommodate space limitations in the main text. MASQUE demonstrates precise makeup application for each prompt, regardless of the targeted region or the subject’s demographics.
To evaluate color variation, we provide examples featuring both a male and a female subject. Additionally, by testing prompts that target different facial regions (eyes, lips, cheeks, skin, and full-face makeup), we demonstrate MASQUE’s ability to generate realistic makeup across various regions. We further provide visualizations without optimization to illustrate its impact.

### v3

@@ 4.2 Detailed Design of MASQUE
To address the limitations of existing AFR methods discussed in Section 4.1, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced. Figure 2 illustrates the pipeline of MASQUE with its pseudocode detailed in Algorithm 1 in Appendix A. The makeup generation process is guided by a user-defined text prompt , refined via cross-attention fusion with a mask. Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR models without compromising visual fidelity. In particular, our method leverages the state-of-the-art Stable Diffusion (Rombach et al., 2022) adapted to generate high-quality adversarial makeup by iterative denoising a random noise latent conditioned on a text embedding . To be more specific, Stable Diffusion is trained to predict the added noise via:
@@
Enhancing Semantic Edits & Locality. To ensure precise localization, we generate a mask that defines the region for modification. To achieve this, we embed the prompt using a Sentence Transformer (Reimers & Gurevych, 2019) model and compare it to embeddings of predefined facial regions. The closest match determines the relevant area for the edit. For instance, if the prompt specifies “a face with red lipstick”, the model identifies lips as the target and generates a lip-area mask. This ensures that our adversarial makeup perturbations remain focused on the correct facial region and do not unintentionally alter other parts of the face. Once the target region is determined, we enhance the influence of makeup-related tokens by maximizing their attention within . This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality (Mao et al., 2023). Specifically, we optimize:
@@
Pairwise Adversarial Guidance. Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios. By contrast, our approach leverages a pair of face images from the same individual, where serves as the guide image. This strategy highlights a significant issue with using only the distance from the original image as the adversarial loss. Diffusion models are designed to reconstruct images similar to the original, so maximizing distance alone creates conflicting objectives. As a result, this approach often produces unstable performance in both image quality and adversarial effectiveness. In our evaluations, we also study the impact of the number of guide images on MASQUE’s performance (Appendix C.1) and whether using self-augmentation without a separate guide image () can be a competitive alternative (Appendix C.2).
Balancing Protection Strength & Image Quality. We introduce adversarial perturbations during the later stage of the diffusion process, ensuring the coarse structure remains intact while subtly altering identity-specific features. To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure:
@@ 5.1 Experimental Setup
To reflect the high-quality nature of online face images, we use high-resolution images () from each of the CelebA-HQ (Karras et al., 2017) and VGG-Face2-HQ (Chen et al., 2024) datasets. For each dataset, we randomly sample 100 identities, each with three images: a probe (to be protected), a reference (as the guide image), and a gallery image (stored in the attacker’s facial recognition model). Note that the VGG-Face2-HQ dataset contains face images under challenging conditions, such as pose variation and natural occlusions like sunglasses and hair.
We compare MASQUE against several baselines: TIP-IM (Yang et al., 2021), AMT-GAN (Hu et al., 2022), C2P (Shamshad et al., 2023), DiffAM (Sun et al., 2024), and FPP (Salar et al., 2025). For C2P, both its impersonating version, C2P (I), and its dodging version, C2P (D), are assessed. We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 (Deng et al., 2019), IRSE50 (Hu et al., 2018), FaceNet (Schroff et al., 2015), and MobileFace (Chen et al., 2018). Without explicitly mentioning, we set in MASQUE (using a single separate guide image of the same individual). For reference-based methods, we selected images from the makeup dataset (Li et al., 2018) used during their pre-training that best matched the given prompt. In addition to public FR models, we evaluate AFR methods against two commercial APIs: Face++111https://www.faceplusplus.com/face-comparing/ and Luxand222https://luxand.cloud/face-api.
For evaluation, we use DSR to measure the protection strength of AFR methods and LPIPS, PSNR, and SSIM (Wang et al., 2004) to assess visual quality, where lower LPIPS and higher PSNR or SSIM indicate better preservation. Without explicit mention, the results are averaged across three makeup prompts, “red lipstick", “blue eyebrow” and “pink eyeshadow”. We also evaluate prompt adherence with metrics such as concentration and hue accuracy. Appendix B provides more details regarding the hyperparameters of MASQUE, external reference implementation, and metric definitions. Additional experiments, ablations, and visualizations are provided in Appendices C–E.
@@ 5.2 Main Results
High Dodging Success. Table 2 presents the DSR for face verification and rank-1 identification in black-box settings using four widely adopted pre-trained FR extractors. For each target model, the others serve as surrogates, with results averaged across three makeup styles. Table 2 demonstrates that MASQUE with consistently outperforms existing baselines across all configurations, achieving an average DSR of for identification and for verification. Besides, even without a guide image, MASQUE still outperforms TIP-IM, the second-best AFR method, by a large margin, highlighting strong protection under the dodging scenario. We also test our method against two commercial FR APIs in verification mode, which assigns similarity scores from to . As proprietary models with unknown training data and parameters, they provide realistic testbeds for our method. Figure 3 shows that our method achieves the lowest similarity scores across both APIs, confirming effectiveness in open- and closed-source settings and reinforcing real-world applicability.
Visual Quality Preservation. MASQUE achieves superior image quality across multiple evaluation metrics, as summarized in Table 3. While TIP-IM attains the highest PSNR and SSIM due to its small perturbation constraint, these pixel-level metrics often fail to reflect perceptual quality (see Figure 1 for qualitative comparison results). In contrast, our approach prioritizes perceptual consistency, balancing content fidelity and visual realism, as demonstrated by its strong LPIPS performance.
| TIP-IM | AMT-GAN | C2P(I) | C2P(D) | DiffAM | FPP | MASQUE () | MASQUE () | |
| LPIPS () | ||||||||
| PSNR () | ||||||||
@@
Localization & Prompt Adherence. Moreover, we examine the localization capability and prompt adherence of MASQUE based on spatial precision and color accuracy. For spatial adherence, we use concentration, which quantifies the proportion of modifications within a binary mask of the target area. Higher scores indicate better localization, while lower scores suggest spillover. For color precision, we evaluate hue accuracy by extracting modified pixels within the mask, converting them to HSV space, and measuring how many fall within the expected hue range. Figure 4 shows that MASQUE consistently achieves the highest concentration scores and competitive hue accuracy across prompts, ensuring precise localization with accurate color application. In contrast, DiffAM struggles with blue eyebrows, likely due to its framework restricting makeup transfer to specific facial regions, limiting its flexibility. In Appendix D.1, we conduct additional experiments using other distance metrics with respect to both in- and out-mask regions, illustrating the strong localization capability of MASQUE. These results confirm that MASQUE achieves strong prompt adherence by maintaining spatial precision while accurately reflecting the intended color attributes.
| None | Resize | Compression | Gaussian | Blur | Fog | De-Makeup | Denoising | |
| Identification | ||||||||
| Verification |
6 Further Analyses
@@
Controllable Makeup Generation. Figure 5 illustrates the strong controllability of MASQUE over text makeup prompts across three specific dimensions: sequential editing, intensity control, and style control. Makeup can be applied progressively to multiple facial regions, with each layer enhancing protection, as shown by decreasing similarity scores. Intensity is adjustable via text prompts (e.g., “vivid”), affecting appearance with minimal impact on effectiveness. Style prompts range from natural (e.g., “tanned skin”) to extreme (e.g., “clown makeup”), with broader coverage offering stronger protection (see Appendix E for additional visualizations of MASQUE-generated images).
@@
Robustness to Image Transformation. We evaluate the robustness of MASQUE using the same CelebA-HQ images and four FR models used in the main experiments, with the prompt “red lipstick”. We then subject the edited images to a variety of transformations, standard post-processing, such as resize and JPEG compression, natural image corruptions, such as Gaussian noise, blur, and fog, and adaptive variations such as DiffAM’s makeup removal and DDPM-based denoising. Table 4 summarizes the averaged DSR across the four considered public FR models, where Table 11 in the appendix provides the full comparison results. MASQUE maintains high dodging success rates across all cases, indicating strong resilience to real-world degradations and adaptive threats. This robustness may arise from the invariance of FR models, which are trained to withstand such transformations, allowing MASQUE’s protection to persist. Also, we test the performance of MASQUE against adversarial purification (Nie et al., 2022), where the results are presented and discussed in Appendix D.4.
@@ 7 Conclusion
We introduced MASQUE, a diffusion-based AFR framework that enables localized adversarial makeup synthesis via text prompts. Unlike prior AFR methods, MASQUE achieves high DSRs while preserving visual quality without relying on external images, reducing the potential risks of targeted misuse. Our experiments demonstrated that MASQUE outperforms existing baselines, with pairwise adversarial guidance playing a central role. Future work may focus on optimizing the efficiency for real-time applications and studying the robustness against adaptively evolving FR models (Fan et al., 2025) to further strengthen MASQUE’s efficacy for facial privacy protection.
Ethics Statement
Our work only uses public datasets and aims to defend against unauthorized or invasive FR systems, so we believe it does not raise any direct ethical concerns. Unlike prior AFR works that are mostly impersonation-driven, our method is distinctive in that it avoids using any external identity’s face image, thereby reducing the risks of targeted abuse of the technology. Nevertheless, our work may raise dual-use concerns. In particular, the same techniques can potentially be misused to evade legitimate or authorized FR systems, such as those used in security or law enforcement contexts. It is important to anticipate and acknowledge these risks, and we encourage future work to consider safeguards or contextual constraints that limit misuse while preserving individual privacy protections.
Reproducibility Statement
To ensure reproducibility, we provide the details of our implementations, including pseudocode in Appendix A and hyperparameter setup in Appendix B. For datasets, we use public image benchmarks, CelebA-HQ and VGG-Face2, which naturally suggest reproducibility. The code for reproducing our experiments is open-sourced at https://github.com/TrustMLRG/MASQUE.
Appendix A Pseudocode of MASQUE
@@ C.1 Number of Images for Adversarial Guidance
| # Guide Images | Iden. DSR | Veri. DSR | Confidence | Similarity | LPIPS |
@@ C.3 Adversarial Strength versus Image Quality
| TIP-IM | AMT-GAN | C2P (I) | C2P (D) | DiffAM | FPP | MASQUE () | |
| LPIPS () | |||||||
| PSNR () | |||||||
@@
| Method | Identification | Verification | ||||||
| MASQUE | ||||||||
| + Resize | ||||||||
| + Compression | ||||||||
| + Gaussian Noise | ||||||||
| + Blur | ||||||||
| + Fog | ||||||||
| + Makeup Removal | ||||||||
| + DDPM Denoising | ||||||||
C.4 Mask and Perturbation Localization
@@ D.1 Localized Editing
To evaluate whether edits are confined to the desired region, images are divided into in-mask and out-mask regions using a binary mask. The corresponding areas in both images are isolated via element-wise multiplication. The areas are then used to normalize similarity metrics proportionally, ensuring fair comparisons. Specifically, we employ DISTS, LPIPS, and PieAPP as evaluation metrics in this experiment: DISTS measures perceptual dissimilarity based on structure and texture, LPIPS uses deep features, and PieAPP reflects human perceptual preferences. Table 7 compares the effectiveness of localized edits, where higher values indicate stronger localization of perturbations. TIP-IM (Yang et al., 2021), as a noise-based method, applies pixel-wise adversarial perturbations uniformly, resulting in the smallest difference for all metrics due to minimal distinction between in-mask and out-mask regions. In contrast, our method introduces significant perturbations in the in-mask region, leading to poorer metrics there, but achieves the best or second-best results in the out-mask region, indicating minimal disruption to untouched areas.
@@ D.4 Robustness against Adversarial Purification
D.5 Computational Analysis
MASQUE does not require fine-tuning, operating directly with a pretrained Stable Diffusion model. In contrast, Clip2Protect fine-tunes per image (~s), and DiffAM performs per-dataset fine-tuning (~min) for each style or identity. At inference, MASQUE runs a one-time iterative optimization (~s), slightly slower than Clip2Protect (~s) and DiffAM (~s). Despite this, MASQUE offers key advantages: it supports arbitrary prompts and identities without retraining, making it more flexible and general. Under typical AFR scenarios, where protection is applied once before sharing, this adaptability outweighs the modest runtime difference.
D.6 Limitations
@@ Appendix E Additional Visualizations
Additional visual results of image-guided makeup transfer are provided in Figure 8 and Figure 9 to accommodate space limitations in the main text. MASQUE demonstrates precise makeup application for each prompt, regardless of the targeted region or the subject’s demographics. To evaluate color variation, we provide examples featuring both a male and a female subject. Additionally, by testing prompts that target different facial regions (eyes, lips, cheeks, skin, and full-face makeup), we demonstrate MASQUE’s ability to generate realistic makeup across various regions. We further provide visualizations without optimization to illustrate its impact.